OpenStudy (anonymous):

Brad bought a $6 binder and several packs of paper that cost $0.60 each. If his total was $13.20, how many packs of paper did brad buy?

OpenStudy (anonymous):

.6p + 6 = 13.2 .6p = 13.2 - 6 =>p = 7.2/ .6 = 12

OpenStudy (anonymous):

13.20 = 6 +.6x -6 -6 7.20 = .6x 7.20/.6=x 12= x

OpenStudy (anonymous):

thanks